satellite73 (satellite73):

show that every unit fraction i.e. a fraction of the form \(\frac{1}{n}\) can be written as the sum of two other unit fractions

OpenStudy (zarkon):

are the two other unit fractions unique?

OpenStudy (anonymous):

good question

OpenStudy (anonymous):

let me see if i can find two different ways to do it for some unit fraction seems kind of likely, not sure

OpenStudy (zarkon):

a trivial way is \[\frac{1}{2n}+\frac{1}{2n}\]

OpenStudy (zarkon):

so i assume they have to be different

OpenStudy (anonymous):

ok cheater, let me rephrase two different unit fractions

myininaya (myininaya):

a unit fraction that looks like 1/[n(n+1)] can be wrriten as 1/n -1/(n+1)

OpenStudy (anonymous):

then you win!

OpenStudy (paxpolaris):

so n can be negative : eg: 3, -4

OpenStudy (anonymous):

positive @myininaya has it

OpenStudy (anonymous):

does this have anything to do with egyptian fractions

OpenStudy (anonymous):

hmm not that i know of

OpenStudy (zarkon):

you need to add a slight restriction on the value of n (if the two other fractions are different)

OpenStudy (anonymous):

reason I say that is because of this... http://en.wikipedia.org/wiki/Unit_fraction

OpenStudy (anonymous):

yes, i see that now i guess \(n>1\)

myininaya (myininaya):

\[n(n+1)=17 \\ n^2+n-17=0 \\ n=\frac{-1 \pm \sqrt{1-4(-17)}}{2(1)}=\frac{-1 \pm \sqrt{69}}{2} \\ \text{ choose } n=\frac{-1 + \sqrt{69}}{2} \\ \text{ so } n+1=\frac{-1 +\sqrt{69}}{2}+1 =\frac{1+\sqrt{69}}{2} \\ \frac{1}{17}=\frac{1}{\frac{-1 +\sqrt{69}}{2}}-\frac{1}{\frac{1+\sqrt{69}}{2}}\] I guess that is still consider a sum (or difference) of unit fractions

myininaya (myininaya):

we did't say n had to be an integer

OpenStudy (anonymous):

you had it at step one

OpenStudy (zarkon):

"A unit fraction is a rational number written as a fraction where the numerator is one and the denominator is a positive integer." ;)

OpenStudy (anonymous):

\[\frac{1}{n(n+1)}=\frac{1}{n}-\frac{1}{n+1}\\ \frac{1}{n}=\frac{1}{n+1}+\frac{1}{n(n+1)}\]

myininaya (myininaya):

oh wow I'm so dumb

myininaya (myininaya):

I could have just solved my thingy for 1/n

myininaya (myininaya):

now you can write 1/17 as a sum of actual unit fractions (now i know the definition of unit fractions which makes sense for that to be the definition)

OpenStudy (anonymous):

i'm so dumb lol

myininaya (myininaya):

no me more let's argue with each other who is more dumb it will be fun

myininaya (myininaya):

another awesome question :)

OpenStudy (anonymous):

no dumb jokes, too easy